OpenStudy (anonymous):

for some reason, i'm getting 101,619.7553 instead of what's in the answer choices, so what am i doing wrong? Mr. Smith wishes to retire in 16 years. When he retires he would like to have $500,000 in his bank account. Mr. Smith's bank pays 10% per year compounded annually. How much should he deposit now to attain his goal? a) $108,816.57 b) $108,811.57 c) $108,813.57 d) $108,814.57 e) $108,810.57

OpenStudy (mathmate):

How did you get 101,619.7553?

OpenStudy (anonymous):

i used the a=P(1+r/n)^nt formula

OpenStudy (anonymous):

same as the last question: to find compound intrest: FutureValue = StartingValue(r/n +)nt FutureValue = future value of the investment StartingValue = Starting deposit r = interest rate n = times annually interest is compunded t = years of the investment total value including interest .

OpenStudy (anonymous):

what are u adding to in this formula? +)nt

OpenStudy (mathmate):

In your case, n=1 (compounded once a year)

OpenStudy (anonymous):

n = times annually interest is compunded t = years of the investment total value including interest .

OpenStudy (mathmate):

So it reduces to A=P(1+r)^t Perhaps you'd like to recheck the calculations.

OpenStudy (anonymous):

n = 1 t = 16 so to find the answer it would be StartingValue(10/1 +)116 just do that will each amount listed in the answer choices to find the correct answer e.g... for a you would do 108816.57(10/1 +)116 if that adds to 500,000 or more that is your answer

OpenStudy (anonymous):

using that formula, answer a will give you $500,009.20 in 16 years.

OpenStudy (anonymous):

thanks, i got the answer now :)

OpenStudy (anonymous):

so, would this problem use the same formula as A=P(1+r)^t? Jack and Tracy would like to have $12,000 for a down payment on a new condo. A savings account at their local credit union will pay 2% per year compounded monthly. How much should they deposit now in this account to have the down payment in 3 years?

OpenStudy (anonymous):

d would give you $500,000.01 since that is only a penny off from the question i would go with that one, on your previous question. hang on ill work on the next question

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

im sorry, i have 2 screaming kids ive gota get off here and deal with, please read over http://www.mathsisfun.com/money/compound-interest.html it probably can explain it better than me, ill be back on in a few minutes if you can figure it out

OpenStudy (anonymous):

reqd amt is 500000/((1+.1)^16)=500000/((1.1)^16) = 108,814.567895 =108,814.57 (appx)

OpenStudy (mathmate):

For the downpayment problem, you need to use the general formula: "i used the a=P(1+r/n)^nt formula" because there are 12 payments per year, so n=12. r=2%=0.02 Amount, a is 12000 (in three years) the time duration, t = 3 years P is the monthly payment. You can transpose the equation to get \( P=\frac{a}{(1+\frac{r}{n})^{nt}} \) so that you can get your answer directly. Feel free to post your answer for a check.